# Capillary levelling of thin liquid films of power-law rheology

### 21 Feb 2024 v1

@@ 3 Comparison with numerical simulations
In order to test the universality of the solutions found in Section 2, we simulate the time-dependent lubrication equation (2). We use a finite volume method, which naturally solves the weak formulation of a conservative transport equation. This property is particularly important in the present case as for $n<1$, we do not expect (2) to have classical (four-times differentiable) solutions. We consider a domain $x\in[0,L]$ divided into $N$ cells of width $\Delta x=L/N$, with the $j$th cell centre at $x_{j}=\Delta x(j-1/2)$, $1<j<N$. The finite difference expression for the third derivative of $h$ on the face between the $j$th and $(j+1)$th cell is
$h_{xxx}\approx\frac{1}{\Delta x^{3}}\left(-h_{j-1}+3h_{j}-3h_{j+1}+h_{j+2}\right),$
where $h_{j}$ is taken to be the representative value of $h$ in cell $j$, while the value of the thickness itself $h$ on the face is given by the average $(h_{j+1}+h_{j})/2$. These expressions are used to approximate the fluxes $q_{j}$. The transport equation (2) then gives
@@
No flux conditions are imposed at the two ends $x=0$ and $x=L$, while no-slope conditions ($h_{x}=0$) are used to define ghost node values that allow the third derivative to be computed on all internal faces. The numerical scheme is then advanced in time using MATLAB’s ode15s implicit time-stepping algorithm.
Using this method we compute solutions of (2) for the two values of the power law exponent we have focussed on thus far ($n=1/2$ and $n=3/2$), and initial conditions that test the behaviour of periodic and localised solutions, respectively. To test the behaviour of periodic solutions, we choose a domain size of $L=2\pi$ and an initial condition
@@
where $\bar{A}$ is the ($n$-dependent) amplitude of the solution to the profile function $\bar{H}(x)$, and $t_{0}$ is the finite levelling time. For $n=1/2$ and $n=3$, then, these become
@@
Figure 4: Numerical simulation of a periodic solution for (a,b) $n=1/2$, and (c,d) $n=3/2$, using the method described in Section 3. (a,c) depict the profile evolution, with initial condition shown as the black dashed curve. In (b) we observe the predicted power-law decay in amplitude for shear-thinning flow, with $A\sim 1.246t^{-1}$ (shown as a black dotted line). In (d) we observe the predicted finite time levelling in the shear-thickening case, with $A\sim 0.044(t_{0}-t)^{3}$ (shown as a black dotted line); here the finite levelling time is approximately $t_{0}\approx 1.8$. Blue circles in (b.d) correspond to the times at which profiles are plotted in (a,c), respectively.
@@ 4 Discussion
The rheology itself may be explicitly regularised, using, for example, a Carreau-type model [Myers2005], which acts like a power law fluid except for small strain rate, where it behaves in a Newtonian manner. For more complex rheology, the flux becomes more complicated to compute, although it is possible to make progress using the method described in [Pritchard2015] (see also [Hinton2022] for the two-dimensional case). Under such a regularisation, a thin film may be expected to level in the same way as described in this paper, until the film becomes very close to level, at which point the Newtonian regime will be reached, and perturbations will switch to exponential decay according to the linear stability analysis of the Newtonian case.
Extending to two spatial dimensions is also of interest. Presumably, there exists a radially symmetric similarity solution that would describe the decay of localised perturbations. For periodic perturbations, one could imagine a radial solution for a thin film in a finite circular domain, or, for a more general shape, one would have to solve the nonlinear elliptic equation that is the counterpart to (6) in higher dimensions. The power law decay of the amplitude in time, however, would be universal.

### 10 Apr 2024 v2

@@ 3 Comparison with numerical simulations
In order to test the universality of the solutions found in Section 2, we simulate the time-dependent lubrication equation (2). We use a finite volume method [Patankar1980], in which fluxes are explicitly computed. This property is particularly important in the present case as for $n<1$, we do not expect (2) to have four-times differentiable solutions, whereas the flux $q$ should be differentiable. We consider a domain $x\in[0,L]$ divided into $N$ cells of width $\Delta x=L/N$, with the $j$th cell centre at $x_{j}=\Delta x(j-1/2)$, $1<j<N$. The finite difference expression for the third derivative of $h$ on the face between the $j$th and $(j+1)$th cell is
$\frac{\partial^{3}h}{\partial x^{3}}\approx\frac{1}{\Delta x^{3}}\left(-h_{j-1}+3h_{j}-3h_{j+1}+h_{j+2}\right),$
where $h_{j}$ is taken to be the representative value of $h$ in cell $j$, while the value of the thickness $h$ itself on the face is given by the average $(h_{j+1}+h_{j})/2$. These expressions are used to approximate the fluxes $q_{j}$. The transport equation (2) then gives
@@
Zero-flux conditions are imposed at the two ends $x=0$ and $x=L$, while zero-slope conditions ($\partial h/\partial x=0$) are used to define ghost node values that allow the third derivative to be computed on all internal faces. The numerical scheme is then advanced in time using MATLAB’s ode15s implicit time-stepping algorithm.
Using this method we compute solutions of (2) for the two values of the power-law exponent we have focussed on thus far ($n=1/2$ and $n=3/2$), and initial conditions that test the behaviour of periodic and localised solutions, respectively. To test the behaviour of periodic solutions, we choose a domain size of $L=2\pi$ and an initial condition
@@
where $\bar{A}$ is the ($n$-dependent) amplitude of the solution to the profile function $\bar{H}(x)$, and $t_{0}$ is the finite levelling time. For $n=1/2$ and $n=3/2$, then, these become
@@
Figure 4: Numerical simulation of a periodic solution for (a,b) $n=1/2$, and (c,d) $n=3/2$, using the method described in Section 3. (a,c) depict the profile evolution for each case, with initial condition shown as the black dashed curve. In (b) we observe the predicted power-law decay in amplitude for shear-thinning flow, with $A\sim 1.246t^{-1}$ (shown as a black dotted line). In (d) we observe the predicted finite time levelling in the shear-thickening case, with $A\sim 0.044(t_{0}-t)^{3}$ (shown as a black dotted line); here the finite levelling time is approximately $t_{0}\approx 1.8$. Blue circles in (b,d) correspond to the times at which profiles are plotted in (a,c), respectively.
@@ 4 Discussion
The rheology itself may be explicitly regularised, using, for example, a Carreau-type model [Myers2005], which acts like a power-law fluid except for small strain rate, where it behaves in a Newtonian manner. For more complex rheology, the flux becomes more complicated to compute, although it is possible to make progress using the method described in [Pritchard2015] (see also [Hinton2022] for the two-dimensional case). Under such a regularisation, a thin film may be expected to level in the same way as described in this paper, until the film becomes very close to level, at which point the Newtonian regime will be reached, and perturbations will switch to exponential decay according to the linear stability analysis of the Newtonian case.
Extending to two spatial dimensions is also of interest. Presumably, there exists a radially symmetric similarity solution that would describe the decay of localised perturbations. For periodic perturbations, one could imagine a radial solution for a thin film in a finite circular domain, or, for a more general shape, one would have to solve the nonlinear elliptic equation that is the counterpart to (7) in higher dimensions. The power-law decay of the amplitude in time, however, would be universal.